How many milliliters of water would you add to 75 milliliters of a 15% acid solution to make a 10% solution? (This is supposed to be solved using a system of linear equations.
Answer by mananth(16946)

------ percent ---------------- quantity
Acid 15 ----------------75 ml
water 0 ----------------x ml
Total 10 ----------------75+x ml

15*75+0*x=10(75+x)

1125=750+10x

-10x=750-1125

-10x=-375

/-10

x=37.5 ml water